# Forced Symmetric Formation Control

### v1

@@ I Introduction
In this paper we will focus on configurations that are constrained by pairwise distances. Such systems are commonly known as (bar-joint) frameworks. Since checking a framework for rigidity is in general very difficult (as it requires solving a system of quadratic equations), a common approach is to check for the linearized (and stronger) notion of rigidity known as ”infinitesimal rigidity.” See Section II-A for precise definitions.
The seminal work from [6] provided the first formal result showing that (minimal) infinitesimal rigidity (MIR) of the interaction network in a team of integrator agents is required to ensure that the gradient controller (locally) converges to the correct formation shape. In the Euclidean space $\mathbb{R}^{2}$, MIR translates to having $32n-$ 3 constraints, where $n$ is the number of agents. Since the work of [6], there has been an explosion of research focusing on the formation control problem from a rigidity theory perspective; see the following for an overview [7, 8, 9, 10, 11]. Within the controls community, the main concern of these works, however, focuses on understanding the resulting dynamics of the control strategies, with most efforts on the stability and convergence properties of these systems [12, 13, 14]. In these works, the assumption of infinitesimal rigidity is taken as a kind of architectural requirement for solving the formation control problem. That is, there has not been a concerted effort to exploit results from rigidity theory to relax the infinitesimal rigidity assumption. In this paper, we will pursue this using recent advances in the rigidity analysis of symmetric frameworks.
@@
This initial work, however, did not leverage the full potential of recent results characterizing symmetry-forced rigidity. In the present article, we extend the results of [15] in the following ways. We develop a concrete mathematical foundation for solving the multi-agent formation control problem under symmetry constraints. The theory provides the symmetric counterpart to the ordinary rigidity-based formation control using the modern language of graph rigidity, and in particular that of forced-symmetric rigidity theory. Notably, we show that $(1+1/|\Gamma|)n$ edges are sufficient when the underlying symmetry group is $\Gamma$. This is significantly smaller than the bound for infinitesimal rigidity. In this direction, we first present a detailed overview of results from the rigidity theory of symmetric frameworks. The central objects in this study are the quotient graphs, providing a graph-theoretic characterization of the vertex and edge orbits induced by a given symmetry group, and the orbit rigidity matrix, which can be thought of as the rigidity matrix associated to the quotient graph of a framework. The orbit rigidity matrix is then used to define an orbit rigidity formation potential for solving the forced-symmetric formation control problem. We then provide a stability and convergence analysis of this control law. Finally, we propose a consensus-augmented version of the forced-symmetric formation control problem that ensures the formation converges to a point other than a globally defined origin. Numerous examples are provided throughout to illustrate the main concepts and results.
@@ II Preliminaries from geometric rigidity
This section provides an overview of basic concepts from geometric rigidity theory and the formation control literature. It will serve as the foundation for the main results in this work.
@@ II-A Rigidity Theory
Since for $d\geq 2$, it is NP-hard to determine if a given framework is rigid, a common approach to study the rigidity of frameworks is to linearise the problem by differentiating the length constraints on the edges. This leads to the notion of infinitesimal (or equivalently, static) rigidity. An infinitesimal motion of a framework $(\mathcal{G},p)$ in $\mathbb{R}^{d}$ is an assignment of velocity vectors, one to each vertex, $u:\mathcal{V}\to\mathbb{R}^{d}$, such that
$\langle p_{i}-p_{j},u_{i}-u_{j}\rangle=0\quad\textrm{ for all }ij\in\mathcal{E}\textrm{,}$ (1)
where $p_{i}=p(i)$ and $u_{i}=u(i)$ for each $i\in\mathcal{V}$. An infinitesimal motion $u$ of $(\mathcal{G},p)$ is trivial if there exists a skew-symmetric matrix $S$ and a vector $c$ such that $u_{i}=Sp_{i}+c$ for all $i\in\mathcal{V}$, and $(\mathcal{G},p)$ is infinitesimally rigid if every infinitesimal motion of $(\mathcal{G},p)$ is trivial, and infinitesimally flexible otherwise.
The ||×$|\mathcal{E}|\times d|\mathcal{V}|$ matrix corresponding to the linear system in (1) is called the rigidity matrix of $(\mathcal{G},p)$, denoted as $R(\mathcal{G},p)$. The row of $R(\mathcal{G},p)$ corresponding to the edge $ij\in\mathcal{E}$ is of the form
$\left(\begin{array}[]{ccccc}0\cdots 0&(p_{i}-p_{j})^{T}&0\cdots 0&(p_{j}-p_{i})^{T}&0\cdots 0\end{array}\right).$
We also employ the useful algebraic representation of the rigidity matrix,
@@ II-B Formation Control
$\dot{p}(t)$ $=-\nabla F_{f}(p(t))$ (5)
$=-R(\mathcal{G},p(t))^{T}\left(R(\mathcal{G},p(t))p(t)-\mathrm{\bf d}^{2}\right),$
satisfies $\underset{t\to\infty}{\lim}\|p_{i}(t)-p_{j}(t)\|=\mathrm{\bf d}_{ij}$ for all $ij\in\mathcal{E}$. Here, $R(\mathcal{G},p(t))$ is the rigidity matrix defined in §II-A. The interested reader may refer to [7, 13] for details on the proof of the above result.
@@ III-A Symmetry in graphs
The combinatorial symmetries of a finite simple graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ are described by its group of automorphisms. An automorphism of $\mathcal{G}$ can be loosely understood as a permutation of $\mathcal{V}$ that maps adjacent vertices to adjacent vertices, and non-adjacent vertices to nonadjacent vertices, and hence preserves all structural properties of $\mathcal{G}$. The formal definition is as follows.
@@
An automorphism of the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a permutation $\psi:\mathcal{V}\to\mathcal{V}$ of its vertex set such that
$\psi(v)\psi(u)\in\mathcal{E}\Leftrightarrow vu\in\mathcal{E}.$
@@ III-C The orbit rigidity matrix
Figure 4: The quotient $\Gamma$-gain graphs of the underlying graphs of the frameworks in Figure 3.
@@ IV-A Symmetry potentials and formation stabilization
$\underset{t\to\infty}{\lim}\|p_{i}(t)-p_{j}(t)\|=\mathrm{\bf d}_{ij}$
and
$\underset{t\to\infty}{\lim}\tau(\gamma)(p_{i}(t))=\underset{t\to\infty}{\lim}p_{\gamma(i)}(t)\quad\textrm{for all }\gamma\in\Gamma,i\in\mathcal{V}.$
@@ IV-B Orbit rigidity potentials and formation stabilization
$u_{i}(t)$ $=\sum_{{ij\in\mathcal{E}(\Gamma_{u})}}(\tau({\gamma_{ij}})p_{j}(t)-p_{i}(t)),$ (20)
for each $u\in\mathcal{V}_{0}$.
@@
$\begin{bmatrix}\dot{p}_{0}(t)\\ \dot{p}_{f}(t)\end{bmatrix}$ $=\begin{bmatrix}-\mathcal{O}^{T}(\mathcal{G}_{0},p_{0}(t))\bigg{(}\mathcal{O}(\mathcal{G}_{0},p_{0}(t))p_{0}(t)-\mathrm{\bf d}_{0}^{2}\bigg{)}\\ 0\end{bmatrix}$ (25)
@@
Here $\mathrm{\bf d}_{0}$ are the distance constraints for the edges in $\mathcal{E}_{0}$. It is interesting to observe the structure of this system. The orbit rigidity matrix in (25) plays the role of the rigidity matrix for (5) in preserving the distances.
@@
$\begin{bmatrix}\dot{\bar{\sigma}}(t)\\ \dot{\bar{q}}(t)\end{bmatrix}$ $=-\begin{bmatrix}\mathcal{O}\mathcal{O}^{T}&\mathcal{O}\bar{E}_{0}(\Gamma)\\ \bar{E}_{0}^{T}(\Gamma)\mathcal{O}^{T}&\bar{E}^{T}(\Gamma)\bar{E}(\Gamma)\end{bmatrix}\begin{bmatrix}\bar{\sigma}(t)\\ \bar{q}(t)\end{bmatrix}$ (35)
@@
Here, $\begin{bmatrix}\bar{E}_{0}(\Gamma)^{T}&\bar{E}_{f}(\Gamma)^{T}\end{bmatrix}^{T}=P\bar{E}(\Gamma)$. We refer to (35) as the orbit error system. For notational simplicity we have dropped the explicit dependence of the orbit rigidity matrix on $\mathcal{G}_{0}$ and $p_{0}(t)$. We also can characterize its equilibria by the set
@@
Consider the error system (35) for a $\tau(\Gamma)$-symmetric framework satisfying Assumption 1. The set of equilibria (45) is asymptotically stable. Furthermore, on the set $\mathcal{X}_{0}$, the control $u(t)=0$ and the configuration vector $p(t)$ converges to a fixed point under the dynamics (25).
@@
The derivative of $V$ along the trajectories of (35) is
@@
Note that for any $\rho>0$ the set $\Psi(\rho)=\{\bar{e}\,:\,V(\bar{e})\leq\rho\}$ such that $\bar{e}(0)\in\Psi(\rho)$ is compact and positively invariant with respect to (35). Therefore, by LaSalle’s Theorem, every solution starting in $\Psi(\rho)$ must approach the largest invariant set where $\dot{V}(\bar{e})=0$, which is precisely the set $\mathcal{X}_{0}$.
Now, observe from (25) that the control can be expressed as
@@
Let $p^{\star}$ be the target formation satisfying conditions (i) and (ii) of Problem 1, and assume that $(\mathcal{G},p^{\star})$ is a $\tau(\Gamma)$-symmetric isostatic framework satisfying Assumption 1. Furthermore, assume that Assumption 2 holds and that the matrices $E(\Gamma_{i})$ are constructed using only a spanning tree subgraph of $\mathcal{G}(\Gamma_{i})$. Then the origin is a locally exponentially stable equilibrium of (35).
@@
which indicates that $V(e(t))$ is negative definite for $e(t)\in\Psi(\rho)\setminus\{0\}$. Thus the exponential stability of the equilibrium $e=0$ in the error system (35) is proved. ∎
In summary, minimally forced-symmetric rigidity provides an architecture for Problem 1 that ensures (local) exponential stability to the desired symmetric formation shape. As in rigidity-based formation control strategies, this result is only local since we must still be concerned with, for example, flip ambiguities of the framework. We also note that the control (25) requires fewer edges than the ordinary rigidity-based control.
@@
In the setting of Theorem 4, the control (25) uses at most $(1+1/|\Gamma|)|{\cal V}|$ edges.
@@
Let ${\cal G}_{0}=({\cal V}_{0},{\cal E}_{0})$ be the quotient graph as used in the description of the control. By the assumption that the vertex set of each orbit induces a tree in the communication graph, the control uses $|{\cal E}_{0}|+(|\Gamma|-1)|{\cal V}_{0}|$ edges. It is known that, if $({\cal G},p^{\ast})$ is a $\tau(\Gamma)$-symmetric isostatic framework, then $|{\cal E}_{0}|\leq 2|{\cal V}_{0}|$, see, e.g., [22, Theorem 62.1.4]. This relation is the symmetric analogue of the well known fact that any isostatic framework has $32|{\cal V}|-$ 3 edges. Hence, the number of edges in the control is bounded by $(|\Gamma|+1)|{\cal V}_{0}|$. By the assumption that $\Gamma$ acts freely on ${\cal V}$, every vertex orbit is of size $|\Gamma|$, and hence $|{\cal V}|=|\Gamma||{\cal V}_{0}|$. Thus, $(|\Gamma|+1)|{\cal V}_{0}|=(1+1/|\Gamma|)|{\cal V}|$. ∎
@@
We now consider the graph on $n=10$ nodes shown in Figure 6(a). This is a $\tau(\Gamma)$-symmetric framework with $\Gamma$ corresponding to the rotational group, where $\tau(\gamma)$ is the rotation matrix of $52\pi/$ 5 radians. Note that this graph has $15$ edges. To solve the formation control problem using the standard approach in (5), we would require an additional $2$ edges to ensure minimal infinitesimal rigidity of the framework (17 total). On the other hand, the control strategy in (25) requires only $9$ edges, with the subgraph shown in Figure 6(b). The quotient graph for this framework is shown in Figure 6(c). Figure 6(d) shows the resulting trajectories when running (25).
@@
(d) Trajectories generated from (25).
(e) Trajectories with centroid consensus term, (52), (46).
Figure 6: Results from Example 6. We consider the $\tau(\Gamma)$-symmetric framework in (a) with the $52\pi/$ 5 rotational symmetry. The framework has two vertex orbits, indicated by the blue and green nodes in (b), along with three edges in the edge orbit (red edges in (b)). For the implementation of (25) we only require a spanning tree subgraph from each $\mathcal{G}(\Gamma_{i})$ and the edges in $\mathcal{E}_{0}$, shown in (b). Figure (c) shows the corresponding quotient gain graph. Figures (d) and (e) show the resulting trajectories without and with the additional consensus term, respectively.
Another interesting byproduct of this strategy is that the symmetries used to define the $\tau(\Gamma)$-symmetric framework are defined with respect to a common inertial frame (see Fig. 6(d)). In the sequel we propose a modification to (45) to relax this point.

### v2

@@ I Introduction
In this paper we will focus on configurations that are constrained by pairwise distances. Such systems are commonly known as (bar-joint) frameworks. Since checking a framework for rigidity is in general very difficult (as it requires solving a system of quadratic equations), a common approach is to check for the linearized (and stronger) notion of rigidity known as ”infinitesimal rigidity” (Section II-A).
The seminal work from [6] provided the first formal result showing that (minimal) infinitesimal rigidity (MIR) of the interaction network in a team of integrator agents is required to ensure that the gradient controller (locally) converges to the correct formation shape. In the Euclidean space $\mathbb{R}^{2}$, MIR translates to having $32n-$ 3 constraints, where $n$ is the number of agents. Since the work of [6], there has been an explosion of research focusing on the formation control problem from a rigidity theory perspective; see the following for an overview [7, 8, 9]. Within the controls community, the main concern of these works, however, focuses on understanding the resulting dynamics of the control strategies, with most efforts on the stability and convergence properties of these systems [10, 11, 12]. In these works, the assumption of infinitesimal rigidity is taken as a kind of architectural requirement for solving the formation control problem. That is, there has not been a concerted effort to exploit results from rigidity theory to relax the infinitesimal rigidity assumption. In this paper, we will pursue this using recent advances in the rigidity analysis of symmetric frameworks.
@@
This initial work, however, did not leverage the full potential of recent results characterizing symmetry-forced rigidity. In the present article, we extend the results of [13] in the following ways. We develop a concrete mathematical foundation for solving the multi-agent formation control problem under symmetry constraints. The theory provides the symmetric counterpart to the ordinary rigidity-based formation control using the modern language of graph rigidity, and in particular that of forced-symmetric rigidity theory. Notably, we show that $(1+1/|\Gamma|)n$ edges are sufficient when the underlying symmetry group is $\Gamma$. This is significantly smaller than the bound for infinitesimal rigidity. This improved bound has direct implications for practical implementation as it leads to a reduction of energy consumption and communication bandwidth. Moreover, this approach can be used to augment any multi-agent coordination problem with explicit symmetry constraints, providing a new conceptual solution to these problems. In this direction, we first present a detailed overview of results from the rigidity theory of symmetric frameworks. The central objects in this study are the quotient graphs, providing a graph-theoretic characterization of the vertex and edge orbits induced by a given symmetry group, and the orbit rigidity matrix, which can be thought of as the rigidity matrix associated to the quotient graph of a framework. The orbit rigidity matrix is then used to define an orbit rigidity formation potential for solving the forced-symmetric formation control problem. We then provide a stability and convergence analysis of this control law. Finally, we propose a consensus-augmented version of the forced-symmetric formation control problem that ensures the formation converges to a point other than a globally defined origin. Numerous examples are provided throughout to illustrate the main concepts and results.
@@ II Preliminaries from geometric rigidity
This section provides an overview of basic concepts from geometric rigidity theory and the formation control literature.
@@ II-A Rigidity Theory
Since for $d\geq 2$, it is NP-hard to determine if a given framework is rigid, a common approach to study the rigidity of frameworks is to linearise the problem by differentiating the length constraints on the edges. This leads to the notion of infinitesimal (or equivalently, static) rigidity. An infinitesimal motion of a framework $(\mathcal{G},p)$ in $\mathbb{R}^{d}$ is an assignment of velocity vectors, one to each vertex, $u:\mathcal{V}\to\mathbb{R}^{d}$, such that $\langle p_{i}-p_{j},u_{i}-u_{j}\rangle=0\quad\textrm{ for all }ij\in\mathcal{E}\textrm{,}$ where $p_{i}=p(i)$ and $u_{i}=u(i)$ for each $i\in\mathcal{V}$. An infinitesimal motion $u$ of $(\mathcal{G},p)$ is trivial if there exists a skew-symmetric matrix $S$ and a vector $c$ such that $u_{i}=Sp_{i}+c$ for all $i\in\mathcal{V}$, and $(\mathcal{G},p)$ is infinitesimally rigid if every infinitesimal motion of $(\mathcal{G},p)$ is trivial, and infinitesimally flexible otherwise.
The ||×$|\mathcal{E}|\times d|\mathcal{V}|$ matrix corresponding to the linear system above is called the rigidity matrix of $(\mathcal{G},p)$, denoted as $R(\mathcal{G},p)$. The row of $R(\mathcal{G},p)$ corresponding to the edge $ij\in\mathcal{E}$ is of the form $\left(\begin{array}[]{ccccc}0\cdots 0&(p_{i}-p_{j})^{T}&0\cdots 0&(p_{j}-p_{i})^{T}&0\cdots 0\end{array}\right).$ We also employ the useful algebraic representation of the rigidity matrix,
@@ II-B Formation Control
$\dot{p}(t)$ $=-R(\mathcal{G},p(t))^{T}\left(R(\mathcal{G},p(t))p(t)-\mathrm{\bf d}^{2}\right),$ (4)
satisfies $\underset{t\to\infty}{\lim}\|p_{i}(t)-p_{j}(t)\|=\mathrm{\bf d}_{ij}$ for all $ij\in\mathcal{E}$. Here, $R(\mathcal{G},p(t))$ is the rigidity matrix defined in §II-A. The interested reader may refer to [7, 10] for more details.
@@ III-A Symmetry in graphs
The combinatorial symmetries of a finite simple graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ are described by its group of automorphisms. An automorphism of $\mathcal{G}$ can be loosely understood as a permutation of $\mathcal{V}$ that maps adjacent vertices to adjacent vertices, and non-adjacent vertices to nonadjacent vertices, and hence preserves all structural properties of $\mathcal{G}$.
@@
An automorphism of the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a permutation $\psi:\mathcal{V}\to\mathcal{V}$ of its vertex set such that $\psi(v)\psi(u)\in\mathcal{E}\Leftrightarrow vu\in\mathcal{E}.$
@@ III-C The orbit rigidity matrix
[…]
$\mathrm{id}$$\psi_{2}$11112222(c)
Figure 4: The quotient $\Gamma$-gain graphs of the graphs of the frameworks in Fig. 3.
@@ IV Forced Symmetric Formation Control
Remark 1.
Assumption 2 is required to ensure all agents within the same vertex orbit eventually exchange information with each other. This is analogous to standard connectivity assumptions in consensus algorithms, of which this is a generalization [31]. Relaxation of this assumption would necessitate a more complicated control architecture including distributed observers to estimate the missing relative state information.
@@ IV-A Symmetry potentials and formation stabilization
$\underset{t\to\infty}{\lim}\|p_{i}(t)-p_{j}(t)\|=\mathrm{\bf d}_{ij}\text{ and}\underset{t\to\infty}{\lim}\tau(\gamma)(p_{i}(t))=\underset{t\to\infty}{\lim}p_{\gamma(i)}(t)$
for all $\gamma\in\Gamma,i\in\mathcal{V}.$
@@ IV-B Orbit rigidity potentials and formation stabilization
$u_{i}(t)$ $=\sum_{{ij\in\mathcal{E}(\Gamma_{v})}}(\tau({\gamma_{ij}})p_{j}(t)-p_{i}(t)),$ (17)
for each $v\in\mathcal{V}_{0}$.
@@
$\begin{bmatrix}\dot{p}_{0}(t)\\ \dot{p}_{f}(t)\end{bmatrix}$ $=\begin{bmatrix}-\mathcal{O}^{T}(\mathcal{G}_{0},p_{0}(t))\bigg{(}\mathcal{O}(\mathcal{G}_{0},p_{0}(t))p_{0}(t)-\mathrm{\bf d}_{0}^{2}\bigg{)}\\ 0\end{bmatrix}$
@@
Here $\mathrm{\bf d}_{0}$ are the distance constraints for the edges in $\mathcal{E}_{0}$. It is interesting to observe the structure of this system. The orbit rigidity matrix in (IV-B) plays the role of the rigidity matrix for (4) in preserving the distances.
@@
$\begin{bmatrix}\dot{\bar{\sigma}}(t)\\ \dot{\bar{q}}(t)\end{bmatrix}$ $=-\begin{bmatrix}\mathcal{O}\mathcal{O}^{T}&\mathcal{O}\bar{E}_{0}(\Gamma)\\ \bar{E}_{0}^{T}(\Gamma)\mathcal{O}^{T}&\bar{E}^{T}(\Gamma)\bar{E}(\Gamma)\end{bmatrix}\begin{bmatrix}\bar{\sigma}(t)\\ \bar{q}(t)\end{bmatrix}$
@@
Here, $\begin{bmatrix}\bar{E}_{0}(\Gamma)^{T}&\bar{E}_{f}(\Gamma)^{T}\end{bmatrix}^{T}=P\bar{E}(\Gamma)$. We refer to (IV-B) as the orbit error system. For notational simplicity we have dropped the explicit dependence of the orbit rigidity matrix on $\mathcal{G}_{0}$ and $p_{0}(t)$. We also can characterize its equilibria by the set
@@
Consider the error system (IV-B) for a $\tau(\Gamma)$-symmetric framework satisfying Assumption 1. The set of equilibria (20) is asymptotically stable. Furthermore, on the set $\mathcal{X}_{0}$, the control $u(t)=0$ and the configuration vector $p(t)$ converges to a fixed point under the dynamics (IV-B).
@@
The derivative of $V$ along the trajectories of (IV-B) is
@@
Note that for any $\rho>0$ the set $\Psi(\rho)=\{\bar{e}\,:\,V(\bar{e})\leq\rho\}$ such that $\bar{e}(0)\in\Psi(\rho)$ is compact and positively invariant with respect to (IV-B). Therefore, by LaSalle’s Theorem, every solution starting in $\Psi(\rho)$ must approach the largest invariant set where $\dot{V}(\bar{e})=0$, which is precisely the set $\mathcal{X}_{0}$.
Now, observe from (IV-B) that the control can be expressed as
@@
Let $p^{\star}$ be the target formation satisfying conditions (i) and (ii) of Problem 1, and assume that $(\mathcal{G},p^{\star})$ is a $\tau(\Gamma)$-symmetric isostatic framework satisfying Assumption 1. Furthermore, assume that Assumption 2 holds and that the matrices $E(\Gamma_{i})$ are constructed using only a spanning tree subgraph of $\mathcal{G}(\Gamma_{i})$. Then the origin is a locally exponentially stable equilibrium of (IV-B).
@@
which indicates that $V(e(t))$ is negative definite for $e(t)\in\Psi(\rho)\setminus\{0\}$. Thus the exponential stability of the equilibrium $e=0$ in the error system (IV-B) is proved. ∎
In summary, minimally forced-symmetric rigidity provides an architecture for Problem 1 that ensures (local) exponential stability to the desired symmetric formation shape. As in rigidity-based formation control strategies, this result is local since we must still be concerned with flip ambiguities of the framework. We also note that the control (IV-B) requires fewer edges than the ordinary rigidity-based control.
@@
In the setting of Theorem 4, the control (IV-B) uses at most $(1+1/|\Gamma|)|{\cal V}|$ edges.
@@
Let ${\cal G}_{0}=({\cal V}_{0},{\cal E}_{0})$ be the quotient graph as used in the description of the control. By the assumption that the vertex set of each orbit induces a tree in the communication graph, the control uses $|{\cal E}_{0}|+(|\Gamma|-1)|{\cal V}_{0}|$ edges. It is known that, if $({\cal G},p^{\ast})$ is a $\tau(\Gamma)$-symmetric isostatic framework, then $|{\cal E}_{0}|\leq 2|{\cal V}_{0}|$, see, e.g., [20, Theorem 62.1.4]. This relation is the symmetric analogue of the well known fact that any isostatic framework has $32|{\cal V}|-$ 3 edges. Hence, the number of edges in the control is bounded by $(|\Gamma|+1)|{\cal V}_{0}|$. By the assumption that $\Gamma$ acts freely on ${\cal V}$, every vertex orbit is of size $|\Gamma|$, and hence $|{\cal V}|=|\Gamma||{\cal V}_{0}|$. Thus, $(|\Gamma|+1)|{\cal V}_{0}|=(1+1/|\Gamma|)|{\cal V}|$. ∎
@@
We now consider the graph on $n=10$ nodes shown in Figure 6(a). This is a $\tau(\Gamma)$-symmetric framework with $\Gamma$ corresponding to the rotational group, where $\tau(\gamma)$ is the rotation matrix of $52\pi/$ 5 radians. Note that this graph has $15$ edges. To solve the formation control problem using the standard approach in (4), we would require an additional $2$ edges to ensure minimal infinitesimal rigidity of the framework (17 total). On the other hand, the control strategy in (IV-B) requires only $9$ edges, with the subgraph shown in Figure 6(b). The quotient graph for this framework is shown in Figure 6(c). Figure 6(d) shows the resulting trajectories when running (IV-B).
Remark 2.
One well-known challenge in formation controllers of the form (4) is to avoid the invariant sub-space of co-linear solutions [6]. While the symmetry-forced formation control (IV-B) is not immune to this problem, it depends greatly on the chosen symmetry. For example, rotational symmetries, as in Example 6, do not have co-linear solutions as an invariant set, while a reflection symmetry only has co-linear invariant solutions that are orthogonal to the mirror axis.
@@
(d) Trajectories generated from (IV-B).
(e) Trajectories with centroid consensus term, (IV-C), (21).
Figure 6: Results from Example 6. We consider the $\tau(\Gamma)$-symmetric framework in (a) with the $52\pi/$ 5 rotational symmetry. The framework has two vertex orbits, indicated by the blue and green nodes in (b), along with three edges in the edge orbit (red edges in (b)). For the implementation of (IV-B) we only require a spanning tree subgraph from each $\mathcal{G}(\Gamma_{i})$ and the edges in $\mathcal{E}_{0}$, shown in (b). Figure (c) shows the corresponding quotient gain graph. Figures (d) and (e) show the resulting trajectories without and with the additional consensus term, respectively.
Of note is that the symmetries used to define the $\tau(\Gamma)$-symmetric framework are defined with respect to a common inertial frame (see Fig. 6(d)). In the sequel we propose a modification to (20) to relax this point.